OpenStudy (kmullis6):

The legs of a right triangle are 5m and 12m long. What is the length of the hypotenuse? Explain how you found your answer.

OpenStudy (kmullis6):

Please help me!

OpenStudy (anonymous):

Use pythagorean theorem a^2+b^2=c^2 5^2+12^2=c^2 can you solve for c?

OpenStudy (kmullis6):

I get confused after I find out what the sum of the legs are. I got 169.

OpenStudy (anonymous):

Thats right, so now you have c^2=169 take the square root of 169 to find c

OpenStudy (kmullis6):

13?

OpenStudy (anonymous):

Yep

OpenStudy (kmullis6):

ok thanks!